# Entanglement of General Subregions in Time-Dependent States

### v1

@@ 1 Introduction
Entanglement entropy (EE) is a powerful probe of the dynamical properties of quantum many-body systems. Its evolution following quantum quenches has yielded deep insight into non-equilibrium dynamics in isolated systems. In $(1+1)$-dimensional conformal field theories (CFTs), the imaginary-time path-integral approach provides a systematic framework for analyzing the growth and saturation of EE after global and local quenches [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]; see also the review [33]. This framework has since been widely applied to studies of entanglement dynamics and thermalization.
@@ 3 EE in global quench model
A global quench describes a process in which the Hamiltonian of the system undergoes a sudden and spatially uniform change across the entire space, causing the system to evolve rapidly from a near-ground-state configuration to a nonequilibrium state. Typically, the system is initially prepared in the ground state of a pre-quench Hamiltonian $H_{0}$, and at time $t_{0}=0$, the Hamiltonian is abruptly changed to $H$. Thereafter, the system evolves unitarily under the new Hamiltonian as behavior is simple in the limits $\eta\to 0$ and $\eta\to 1$. When $\eta\to 1$, the bulk identity block dominates and $F(1)=1.$ For $\eta\ll 1$, the short-distance bulk–boundary OPE yields $F(\eta)\simeq(A_{b}^{\Phi})^{2}\,\eta^{x_{b}},$ where $A_{b}^{\Phi}$ is the bulk-to-boundary OPE coefficient and $x_{b}$ is the scaling dimension of the leading boundary operator. For twist operators, the leading boundary operator is the identity, so $x_{b}=0$, and the expression simplifies to $F(\eta)\simeq(A_{b}^{\Phi})^{2}$. For further details, see Ref. [33].
@@ 6 Conclusion
In this work, we have developed a systematic framework for computing entanglement entropy (EE) in $(1+1)$-dimensional conformal field theories (CFTs) for general spacetime intervals, with a particular focus on timelike intervals. By introducing the spacetime density matrix formalism, we have generalized the conventional density matrix to incorporate timelike correlations, ensuring that EE for both spacelike and timelike intervals is well-defined and computable, even in time-dependent states. This framework, utilizing the Schwinger-Keldysh path integral, extends the imaginary-time path-integral approach, enabling a unified treatment of static and dynamical systems.

### v2

@@ 1 Introduction
EE is a powerful probe of the dynamical properties of quantum many-body systems. Its evolution following quantum quenches has yielded deep insight into non-equilibrium dynamics in isolated systems. In $(1+1)$-dimensional conformal field theories (CFTs), the imaginary-time path-integral approach provides a systematic framework for analyzing the growth and saturation of EE after global and local quenches [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]; see also the review [33]. This framework has since been widely applied to studies of entanglement dynamics and thermalization.
@@ 3 EE in global quench model
A global quench describes a process in which the Hamiltonian of the system undergoes a sudden and spatially uniform change across the entire space, causing the system to evolve rapidly from a near-ground-state configuration to a nonequilibrium state. Typically, the system is initially prepared in the ground state of a pre-quench Hamiltonian $H_{0}$, and at time $t_{0}=0$, the Hamiltonian is abruptly changed to $H$. Thereafter, the system evolves unitarily under the new Hamiltonian as
$|\Psi(t_{0})\rangle=e^{-iHt_{0}}|\Psi_{0}\rangle.$ (11)
In CFTs, Calabrese and Cardy proposed an analytical model in which the initial state $|\Psi_{0}\rangle$ is taken to be a conformally invariant boundary state $|B\rangle$, which then evolves over a finite imaginary time $\tau^{\prime}$.
$|\Psi_{0}\rangle\propto e^{-\tau^{\prime}H}|B\rangle.$ (12)
Here, the parameter $\tau^{\prime}$ characterizes the finite correlation length of the initial state, ensuring the regularity of the path integral in the short-distance limit. In this construction, the dynamics of the system can be studied by analytically continuing the Euclidean path integral to real time.
After obtaining the form of the initial state, we can further examine the dynamical evolution of local operators in this quenched background. In CFT, the dynamical information of the system is typically encoded in the correlation functions of local operators. In particular, for the global quench setup, the two-point correlation function
$\langle\Psi_{0}(t)|{\Phi}(x_{0},0){\Phi}(x_{1},t_{1}-t_{0})|\Psi_{0}(t)\rangle$
$=$ $\langle B|e^{-\tau^{\prime}H}e^{iHt_{0}}{\Phi}(x_{0},0){\Phi}(x_{1},t_{1}-t_{0})e^{-iHt_{0}}e^{-\tau^{\prime}H}|B\rangle$
$=$ $\langle B|{\Phi}(x_{0},\tau_{0}){\Phi}(x_{1},\tau_{1})|B\rangle,$ (13)
where the analytic continuation to real time is taken as $\tau_{i}\to\tau^{\prime}+it_{i}$. The above two-point correlation function is defined on a strip geometry with Euclidean width $2\tau^{\prime}$, where both boundaries are equipped with the same conformally invariant boundary condition encoded by the boundary state $|B\rangle$. This strip can be conformally mapped to the upper half-plane (UHP) through the transformation
$w(z)=\frac{2\tau^{\prime}}{\pi}\ln z,$ (14)
which allows correlation functions on the strip to be represented in terms of standard CFT correlators on the UHP.
The above correlator on the strip follows from mapping the upper half-plane (UHP) to the strip and applying Cardy’s general boundary CFT expression for two-point functions of primary operators. Its extension to twist fields for computing EE was developed in the seminal work of Calabrese and Cardy [61] [62]
Therefore, the two-point function on the strip can be written as
$\langle\Phi(w_{0})\Phi(w_{1})\rangle_{\mathrm{strip}}=\left|\frac{\partial w_{0}}{\partial z_{0}}\right|^{-2h_{n}}\left|\frac{\partial w_{1}}{\partial z_{1}}\right|^{-2h_{n}}\left(\frac{z_{0\bar{1}}\,z_{1\bar{0}}}{z_{01}\,z_{\bar{0}\bar{1}}\,z_{0\bar{1}}\,z_{1\bar{1}}}\right)^{2h_{n}}F(\eta),$ (15)
where $\bar{z}_{i}$ denotes the complex conjugate of $z_{i}$, and $z_{i\bar{j}}=z_{i}-\bar{z}_{j}$. The function $F(\eta)$ depends on the cross ratio $\eta=\frac{z_{0\bar{0}}\,z_{1\bar{1}}}{z_{0\bar{1}}\,z_{1\bar{0}}}.$ In general, $F(\eta)$ has no closed-form expression, but its asymptotic behavior is simple in the limits $\eta\to 0$ and $\eta\to 1$. When $\eta\to 1$, the bulk identity block dominates and $F(1)=1.$ For $\eta\ll 1$, the short-distance bulk–boundary OPE yields $F(\eta)\simeq(A_{b}^{\Phi})^{2}\,\eta^{x_{b}},$ where $A_{b}^{\Phi}$ is the bulk-to-boundary OPE coefficient and $x_{b}$ is the scaling dimension of the leading boundary operator. For twist operators, the leading boundary operator is the identity, so $x_{b}=0$, and the expression simplifies to $F(\eta)\simeq(A_{b}^{\Phi})^{2}$. For further details, see Ref. [33].
@@ 6 Conclusion
In this work, we have developed a systematic framework for computing EE in $(1+1)$-dimensional CFTs for general spacetime intervals, with a particular focus on timelike intervals. By introducing the spacetime density matrix formalism, we have generalized the conventional density matrix to incorporate timelike correlations, ensuring that EE for both spacelike and timelike intervals is well-defined and computable, even in time-dependent states. This framework, utilizing the Schwinger-Keldysh path integral, extends the imaginary-time path-integral approach, enabling a unified treatment of static and dynamical systems.